# A practical guide to Digital Micro-mirror Devices (DMDs) for wavefront shaping

### v4

@@ 3.1 Presentation of the problem
While only capable of providing hardware binary amplitude modulation, DMDs serve as a potent tool for wavefront shaping and sensing. These applications critically require characterization and correction of aberrations caused by the non-flatness of the DMD surface. For LC-SLMs, the manufacturer typically characterizes the surface’s inhomogeneities within the plane of the modulator and provides a spatial phase profile of the introduced aberrations based on the operational wavelength. It is noteworthy that when utilized for intensity modulation on a plane conjugated to the DMD plane, as it is done in digital projectors, the system becomes insensitive to the aberrations caused by the DMD surface. Consequently, these effects are commonly overlooked and rarely documented in the information provided by the manufacturers. Nonetheless, they are non-negligible, with deviations from a flat surface that typically spans several wavelengths [24].
@@ 3.2 Finding the correction pattern
In the literature, various methods have been proposed to characterize the phase pattern of the DMD aberrations in the plane of the modulator. Typically, this involves using a model for the aberrations, tweaking the parameters to align with the measurements [25, 14, 24]. An alternative approach entails direct measurement of the distorted wavefront, either employing a wavefront sensor such as a Shack-Hartmann [26], or via interferometry. While such methods yield accurate results, they necessitate adaptation to the particular setup conditions, and frequently require supplementary optical components, meticulous alignment, and custom software. Furthermore, they also demand a meticulously calibrated and stable setup. In this section, we introduce a straightforward method to characterize the aberrations using a lens and a camera. This technique can be employed for any system that offers phase modulation, such as LC-SLMs or deformable mirrors.
@@
Figure 12: Setup for aberration correction. (a) A schematic representation of the DMD array divided into macropixels, of which the phase can be controlled. The precise shape of the macropixels and the area controlled are not crucial for the characterization of aberrations to be effective, provided that the spatial sampling is sufficient to capture the highest spatial frequencies of the aberrations. (b) A schematic representation of the setup employed to characterize the aberrations. A camera is used to image the far field on the DMD surface using a lens. (c) shows the ideal intensity pattern corresponding to the numerical aperture of the illumination setup, i.e. its theoretical ideal point spread function. (d) show the actual recorded intensity pattern corresponding to a distorted point spread function due to the DMD aberrations.
@@
We assume the DMD is configured to deliver phase modulation [11, 12]. This implies that the modulator can be divided into $N$ sections, which we designate as macropixels, where the phase can be controlled independently. We use a lens and a camera in its Fourier plane, illuminating the modulator with a collimated beam that extends over the entire area of the modulator intended for use. In scenarios where there are minimal or no aberrations, the intensity pattern observed would mimic the PSF of the lens, such as an Airy disk depicted in Fig. 12.c. However, in practice, we encounter a substantially distorted pattern, like the one represented in Fig. 12.d. A more detailed depiction of the setup for aberration characterization, in the context of a wavefront shaping application in complex media is presented in Fig. 13.a. In the case where the medium is a multimode fiber, one could leverage the reflection from the input surface to directly visualize the intensity pattern of the input plane, as shown in Fig. 13.b. Note that for the later approach to be reliable, it is necessary to properly align the system to ensure that the plane of the input facet corresponds exactly to the Fourier plane of the DMD, and that the input facet plane is conjugated with the camera plane.
@@
Figure 14: Correction phase map. (a) Schematic representation of the function we want to optimize, i.e. we optimize the ratio between the sum of the intensity of the measured image in a diffraction-limited disk and the sum of the intensity elsewhere. We both want to maximize the intensity of the main lobe of the point spread function and minimize its side lobes. (b) Values of the experimentally obtained optimal coefficients in the Zernike polynomials basis. (c) Resulting optimal phase pattern in the chosen illumination area.
We perform a sequential optimization of parameters $a_{n}$ to maximize a specific function designed to be maximal for the ideal correction of optical aberrations. We first generate a mask, represented by a disk, centered around the point of maximum intensity of the original image (refer to Fig. 12.d) with a radius equivalent to a single speckle grain. The exact size of this radius for a successful optimization is not critical and can be determined by approximating the dimensions of the ideal point spread function, expressed as $r_{0}\approx M\frac{\lambda}{2NA}$, where $NA$ is the numerical aperture of the optical system and $M$ refers to its magnification. For a given output intensity pattern, we compute an element-wise product between this image and the created mask, followed by a summation. This calculated sum is then divided by an analogous product, but with the complementary mask substituted in place of the original in order to minimize the side lobes of the point spread function. We set initial parameter values as $a^{\prime}_{n}=0\,\,\forall n\in[0..N]$. For each parameter, we test different values of $a^{\prime}_{n}$, construct the phases for every micropixel according to $\phi_{i}^{\text{corr}}=\sum_{n=0}^{N}a^{\prime}_{n}Z_{n}(r_{i},\theta_{i})$, record the resulting intensity profile, and evaluate the corresponding cost function. For each parameter, the value that results in the maximum output is retained. To mitigate potential noise or instability, this complete process is reiterated $3$ times for each parameter.
To estimate the quality of the correction, we compute the Strehl ratio of the point spread function. It is defined as the maximum of the measured PSF divided by the maximum of the ideal one. This optimal PSF is the squared modulus of the Fourier transform of a circular aperture [28]:
@@
Figure 15: Effect of Zernike polynomials on the optimization of the PSF. (a) Strehl ratio of the point spread function as a function of the number of Zernike polynomials used for the compensation of the aberrations. (b) Intensity profile of the point spread function for different number of Zernike polynomials in linear and logarithmic scale. To generate this data, we take the final solution of the optimization experimental procedure, i.e. the optimal coefficients $a^{\prime}_{n}$, and generate the corrected PSF with an increasing number of Zernike polynomials.
@@ 3.3 Python code example
DMDs are not flat and can introduce aberrations; these are typically much stronger than those commonly observed with liquid crystal SLMs. This can be counteracted in situ using a standard setup and a straightforward optimization procedure to maximize the intensity at the central position of the point spread function.
@@ 4.1 Mechanical stability
A simple yet effective solution consists in dampening the vibrations at the flat cable’s level by clamping it with a soft material, as depicted in Fig. 18. This can be achieved using commonly available materials. In this context, we utilize simple foam, typically used for packaging, and secure it to the cable with two metallic plates, screws, and nuts. We observe a significant decrease in the phase fluctuations, as demonstrated in Fig. 17 (blue curve).
Figure 18: Dumping mechanical vibrations. Pictures of experimental setups where a damping of the vibrations is implemented. (a) The flat cable is secured with foam and metallic plates. (b) The flat cable is in tension against a foam material supported by a metallic plate.
@@ 4.2 Thermal stability
with $\bar{I}(\vec{r},t)=I(\vec{r},t)-\left\langle I(\vec{r},t)\right\rangle_{t}$ and $\left\langle.\right\rangle_{\vec{r}}$. $\left\langle.\right\rangle_{t}$ represent the spatial averaging over the region of interest of the output plane (i.e. the plane of the camera sensor) and the temporal averaging over the measured frames. Fig. 19 shows the measured decorrelation over time when running an sequence in an infinite loop. $t=0$ correspond in the first case (a) to the start of a sequence and is set in the second one to 4.2 hours post the after of the sequence (b). For the scenario (a), it is to note that before the sequence commenced, the DMD was active but remained in the idle state, meaning that it was not executing a sequence. We observe a non-negligible decorrelation over time after the sequence started that slows down after about 1 hour. When running the same experiment 4.2 hours after starting the sequence, the correlation is now stable.
@@
DMDs take about an hour to thermally stabilize when a sequence is running. It can countered by using a thermoelectric cooler and a temperature sensor. It can also be simply mitigated by letting a sequence run for few hours before starting the experiment.
@@ 5 Conclusion
DMDs are powerful tools for wavefront shaping applications in complex media due to their high pixel count, relative low cost, and high refresh rate. However, mostly due to their original purpose of amplitude modulation of incoherent light for video projection, several effects need to be taken into account when using DMDs for wavefront shaping. First, the choice of its pixel pitch must be made carefully by considering the wavelength of operation to ensure a good modulation quality and diffraction efficiency. Furhtermore, the DMD surface is not flat and can introduce aberrations which can be corrected using a simple optimization procedure. Finally, the DMD is sensitive to mechanical vibrations and thermal variations, which can be mitigated ensuring a good mechanical and thermal isolation of the device.

### v5

@@ 3.1 Presentation of the problem
While only capable of providing hardware binary amplitude modulation, DMDs serve as a potent tool for wavefront shaping and sensing. These applications critically require characterization and correction of aberrations caused by the non-flatness of the DMD surface. For LC-SLMs, the manufacturer typically characterizes the surface’s inhomogeneities within the plane of the modulator and provides a spatial phase profile of the introduced aberrations based on the operational wavelength. It is noteworthy that when utilized for intensity modulation on a plane conjugated to the DMD plane, as it is done in digital projectors, the system becomes insensitive to the aberrations caused by the DMD surface. Consequently, these effects are commonly overlooked and rarely documented in the information provided by the manufacturers. However, studies suggest that the residual stresses developed in the thin films during the manufacturing process of the DMDs induce non-negligible surface deformations [24, 25, 26].
@@ 3.2 Finding the correction pattern
In the literature, various methods have been proposed to characterize the phase pattern of the DMD aberrations in the plane of the modulator. Typically, this involves using a model for the aberrations, tweaking the parameters to align with the measurements [27, 12, 26]. An alternative approach entails direct measurement of the distorted wavefront, either employing a wavefront sensor such as a Shack-Hartmann [28], or via interferometry. While such methods yield accurate results, they necessitate adaptation to the particular setup conditions, and frequently require supplementary optical components, meticulous alignment, and custom software. Moreover, once the phase correction map has been obtained, it must then be aligned with the pixel array in order to implement the phase correction on the DMD and compensate for aberrations. In this section, we introduce a straightforward method for characterizing aberrations using a lens and a camera. The goal is to in situ compensate for the aberrations using a phase modulation scheme (Lee holograms), by iteratively optimizing the point spread function (PSF) at the focal plane of a lens. The resulting phase correction map is the conjugate of the one induced by the surface deformation of the DMD, and can subsequently be used to correct aberrations with any desired complex modulation scheme. This technique can be employed for any system that offers phase modulation, such as LC-SLMs or deformable mirrors.
@@
Figure 13: Setup for aberration correction. (a) A schematic representation of the DMD array divided into macropixels, of which the phase can be controlled. The precise shape of the macropixels and the area controlled are not crucial for the characterization of aberrations to be effective, provided that the spatial sampling is sufficient to capture the highest spatial frequencies of the aberrations. (b) A schematic representation of the setup employed to characterize the aberrations. A camera is used to image the far field on the DMD surface using a lens. (c) shows the ideal intensity pattern corresponding to the numerical aperture of the illumination setup, i.e. its theoretical ideal PSF. (d) show the actual recorded intensity pattern corresponding to a distorted PSF due to the DMD aberrations.
@@
We assume that the DMD is configured to deliver phase modulation [15, 16]. It is classically done using Lee holograms or similar approaches, where the optical phase is encoded in the local spatial displacement of a regular array of amplitude fringes displayed on the DMD. By filtering a single diffraction order, the resulting field retains only the phase modulation. Specific implementations and variants of such approaches are discussed in [16]. This implies that the modulator can be divided into $N$ sections, which we designate as macropixels, where the phase can be controlled independently. We use a lens and a camera in its Fourier plane, illuminating the modulator with a collimated beam that extends over the entire area of the modulator intended for use. In scenarios where there are minimal or no aberrations, the intensity pattern observed would mimic the PSF of the lens, such as an Airy disk depicted in Fig. 13.c. However, in practice, we encounter a substantially distorted pattern, like the one represented in Fig. 13.d. A more detailed depiction of the setup for aberration characterization, in the context of a wavefront shaping application in complex media is presented in Fig. 14.a. In the case where the medium is a multimode fiber, one could leverage the reflection from the input surface to directly visualize the intensity pattern of the input plane, as shown in Fig. 14.b. Note that for the later approach to be reliable, it is necessary to properly align the system to ensure that the plane of the input facet corresponds exactly to the Fourier plane of the DMD, and that the input facet plane is conjugated with the camera plane.
@@
Figure 15: Correction phase map. (a) Schematic representation of the function we want to optimize, i.e. we optimize the ratio between the sum of the intensity of the measured image in a diffraction-limited disk and the sum of the intensity elsewhere. We both want to maximize the intensity of the main lobe of the PSF and minimize its side lobes. (b) Values of the experimentally obtained optimal coefficients in the Zernike polynomials basis. (c) Resulting optimal phase pattern in the chosen illumination area.
We perform a sequential optimization of parameters $a_{n}$ to maximize a specific function designed to be maximal for the ideal correction of optical aberrations. The approach is inspired by methods used for focusing through complex disordered media [1], by effectively treating the aberrations of the DMD as equivalent to propagation through an unknown linear medium. For each degree of freedom (i.e. the phase on one group of pixels in Ref. [1], or the coefficient of one Zernike polynomial in this section), we scan different values and retain the one that maximizes the intensity at the target location. We first generate a mask, represented by a disk, centered around the point of maximum intensity of the original image (refer to Fig. 13.d) with a radius equivalent to a single speckle grain. The exact size of this radius for a successful optimization is not critical and can be determined by approximating the dimensions of the ideal PSF, expressed as $r_{0}\approx M\frac{\lambda}{2NA}$, where $NA$ is the numerical aperture of the optical system and $M$ refers to its magnification. For a given output intensity pattern, we compute an element-wise product between this image and the created mask, followed by a summation. This calculated sum is then divided by an analogous product, but with the complementary mask substituted in place of the original in order to minimize the side lobes of the PSF. We set initial parameter values as $a^{\prime}_{n}=0\,\,\forall n\in[0..N]$. For each parameter, we test different values of $a^{\prime}_{n}$, construct the phases for every micropixel according to $\phi_{i}^{\text{corr}}=\sum_{n=0}^{N}a^{\prime}_{n}Z_{n}(r_{i},\theta_{i})$, record the resulting intensity profile, and evaluate the corresponding cost function. For each parameter, the value that results in the maximum output is retained. To mitigate potential noise or instability, this complete process is reiterated $3$ times for each parameter.
To estimate the quality of the correction, we compute the Strehl ratio of the PSF. It is defined as the maximum of the measured PSF divided by the maximum of the ideal one. This optimal PSF is the squared modulus of the Fourier transform of a circular aperture [30]:
@@
Figure 16: Effect of Zernike polynomials on the optimization of the PSF. (a) Strehl ratio of the PSF as a function of the number of Zernike polynomials used for the compensation of the aberrations. (b) Intensity profile of the PSF for different number of Zernike polynomials in linear and logarithmic scale. To generate this data, we take the final solution of the optimization experimental procedure, i.e. the optimal coefficients $a^{\prime}_{n}$, and generate the corrected PSF with an increasing number of Zernike polynomials.
@@ 3.3 Python code example
DMDs are not flat and can introduce aberrations; these are typically much stronger than those commonly observed with liquid crystal SLMs. This can be counteracted in situ using a standard setup and a straightforward optimization procedure to maximize the intensity at the central position of the PSF.
@@ 4.1 Mechanical stability
A simple yet effective solution consists in damping the vibrations at the flat cable’s level by clamping it with a soft material, as depicted in Fig. 19. This can be achieved using commonly available materials. In this context, we utilize simple foam, typically used for packaging, and secure it to the cable with two metallic plates, screws, and nuts. We observe a significant decrease in the phase fluctuations, as demonstrated in Fig. 18 (blue curve).
Figure 19: Damping mechanical vibrations. Pictures of experimental setups where a damping of the vibrations is implemented. (a) The flat cable is secured with foam and metallic plates. (b) The flat cable is in tension against a foam material supported by a metallic plate.
@@ 4.2 Thermal stability
with $\bar{I}(\vec{r},t)=I(\vec{r},t)-\left\langle I(\vec{r},t)\right\rangle_{t}$ and $\left\langle.\right\rangle_{\vec{r}}$. $\left\langle.\right\rangle_{t}$ represent the spatial averaging over the region of interest of the output plane (i.e. the plane of the camera sensor) and the temporal averaging over the measured frames. Figure 20 shows the measured decorrelation over time when running an sequence in an infinite loop. $t=0$ correspond in the first case (a) to the start of a sequence and in the second case (b) is set to 4.2 hours after the start of the sequence. For the scenario (a), it is worth noting that before the sequence commenced, the DMD was active but remained in the idle state, meaning that it was not executing a sequence. We observe a non-negligible decorrelation over time after the sequence started that slows down after about 1 hour. When running the same experiment 4.2 hours after starting the sequence, the correlation is now stable.
@@
DMDs take about an hour to thermally stabilize when a sequence is running. It can be countered by using a thermoelectric cooler and a temperature sensor. It can also be simply mitigated by letting a sequence run for few hours before starting the experiment.
@@ 5 Conclusion
DMDs are powerful tools for wavefront shaping applications in complex media due to their high pixel count, relative low cost, and high refresh rate. However, mostly due to their original purpose of amplitude modulation of incoherent light for video projection, several effects need to be taken into account when using DMDs for wavefront shaping. First, the choice of its pixel pitch must be made carefully by considering the wavelength of operation to ensure a good modulation quality and diffraction efficiency. Furthermore, the DMD surface is not flat and can introduce aberrations which can be corrected using a simple optimization procedure. Finally, the DMD is sensitive to mechanical vibrations and thermal variations, which can be mitigated ensuring a good mechanical and thermal isolation of the device.
@@
Appendix A: 1D calculation of the DMD diffraction effect
Assuming the effect of the device’s finite size and illumination to be negligible, we can represent the field reflected from the device under the influence of plane wave illumination across two systems as follows. Here, $\theta_{B}$ denotes the blaze angle, i.e. the tilt of the micromirror normal with respect to the device normal, so that the corresponding specular direction is $2\theta_{B}-\alpha$:
$R_{\text{flat}}(x)$ $\propto\left[\Pi\left(x/d^{\prime}\right)\otimes_{x}\sum_{k}\delta(x-kd)\right]e^{j\frac{2\pi}{\lambda}\sin(\alpha)x}\,$, (11)
$R_{\text{blazed}}(x)$ $\propto\underbrace{\left(\underbrace{\Pi\left(x/d^{\prime}\right)}_{\text{pixel size}}\,\underbrace{e^{j\frac{2\pi}{\lambda}\left(\sin(2\theta_{B}-\alpha)-\sin(\alpha)\right)x}}_{\begin{subarray}{c}\text{blazed angle}\,,\\ \text{+ angle of incidence}\end{subarray}}\right)}_{\text{pixel response}}\otimes_{x}\underbrace{\sum_{k}\delta(x-kd)}_{\text{periodicity}}\,e^{j\frac{2\pi}{\lambda}\sin(\alpha)x}\,.$
with $\Pi(x)$ the rectangular function, representing the finite size of the pixel, defined as:
$\Pi(x)=\begin{cases}1,&\text{if }-\frac{1}{2}<x<\frac{1}{2},\\ 0,&\text{otherwise}.\end{cases}$ (12)
The filling fraction, in the 1D case, is the ratio of the size of the pixels (or micro-mirrors) $d^{\prime}$, and the pitch $d$, such that $d^{\prime}=\rho d$.
The Fourier transform of Eq. 11 can be written as:
$I_{\text{flat}}(\theta)\propto\sum_{p}\delta\bigl(\sin(\theta)+\sin(\alpha)-p\,\sin(\theta_{D})\bigr)\times\operatorname{sinc}^{2}\left(\pi\rho\frac{\sin(\theta)-\sin(\alpha)}{\sin(\theta_{D})}\right)$ (13)
$I_{\text{blazed}}(\theta)\propto\underbrace{\sum_{p}\delta\bigl(\sin(\theta)+\sin(\alpha)-p\,\sin(\theta_{D})\bigr)}_{\text{orders of diffraction}}\times\underbrace{\operatorname{sinc}^{2}\left(\pi\rho\frac{\sin(\theta)-\sin(2\theta_{B}-\alpha)}{\sin(\theta_{D})}\right)}_{\text{envelope}}\,.$
We observe that the envelope (right-hand term) is maximal for $\sin(\theta_{\text{max}})=\sin(2\theta_{B}-\alpha)$, while the effect of the periodicity (left-hand term) is maximal when $\sin(\theta_{p})+\sin(\alpha)=p\,\lambda/d$, representing the orders of diffraction.
Appendix B: 2D calculation of blazed grating condition
To analyze more precisely the effect of diffraction in a DMD, one needs to consider the 2D surface of the modulator. We can establish a Cartesian coordinate system on the plane of the DMD, with axes $x$ and $y$ aligned with the pixel sides (refer to Fig. 5a). The pixels are uniformly repeated along these axes. However, a technical challenge arises in that the axis of rotation of the pixels aligns with the pixel diagonals, resulting in a rotation by 45 degrees with respect to the $x$ and $y$ axes. For the convenience of alignment and manipulation of the optical setup, it is preferable to work with the incident and outgoing beams which have the optical axis contained in the horizontal plane, i.e. a plane parallel to the table surface. A straightforward and common solution is to rotate the chip by 45 degrees relative to the horizontal plane, which aligns the pixel axis of rotation to be vertical.
with $\vec{k}_{in}$ the incoming wavevector and $\vec{r}$ the position in the modulator plane, represented by the coordinate $(x,y)$.
The total phase then reads:
$\phi(x,y)=\phi(x+d,y)=\phi(x,y+d)\,.$ (17)
The new blazed grating equation then reads :